A particle is projected in vertical plane such that after 5 seconds its velocity becomes horizontal. The maximum height attained by the projectile from point of projection is n3 m. The value of n is _______. (g = 10 m/s2)

Grade:11

1 Answers

Forum 2
one year ago
Hello student

Velocity of projectile=62+82​=100​=10ms−1
Angle of projectileθ=tan−168​=53∘
Range of projectile is given byR=gu2sin2θ​
R=10100×sin106∘​=9.6m
hope you find the answer helpful